Albert wants to enlarge his living room of (x^2 + 10x + 24) square feet to extend into the backyard. The length of the living room is less than the width.
After construction, the area will be (x^2 + 14x + 45) square feet.
a. Find the dimensions of the living room before construction.
b. Find the dimensions of the living room after construction.
**pls show ur work so i can understand how to do this question, thank you!

Answer by ikleyn(52787)

.

It was  x^2 + 10x + 24 = factor left side = (x+6)*(x+4).

Original dimensions were (x+6) feet by (x+4) feet.



It became  x^2 + 14x + 45 = factor left side = (x+9)*(x+5).

Final dimensions became (x+9) feet by (x+5) feet.


Solved, with explanations. The work is shown in full.